OpenStudy (anonymous):

Can someone please help me find the center, vertices, foci, eccentricity, lengths of minor and major axis of \[\frac{ (x-5)^2 }{ 4 }+\frac{ (y+3)^2 }{ 16 }=1\]

OpenStudy (anonymous):

oh no not again ...

OpenStudy (anonymous):

thats what i said :( last problem though so thats my motivation! :p

OpenStudy (anonymous):

you know what this is right?

OpenStudy (anonymous):

is it a parabola?

OpenStudy (anonymous):

oh no an ellips!

OpenStudy (anonymous):

yes it is do you know what it looks like (we need that to start)

OpenStudy (anonymous):

 looks like that :)

OpenStudy (anonymous):

wow that is ugly but yes you got the center?

OpenStudy (anonymous):

sorry lol, i think the center is (5,-3) but im not sure

OpenStudy (anonymous):

yes you are right (not sure why you were not sure)

OpenStudy (anonymous):

im just never sure of myself in math :(

OpenStudy (anonymous):

i guess that is good ok next for the vertices which is very much like the last one

OpenStudy (anonymous):

\[\frac{ (x-5)^2 }{ 4 }+\frac{ (y+3)^2 }{ 16 }=1\]in this case since \(16>4\) we have \[\frac{ (x-h)^2 }{ b^2 }+\frac{ (y-k)^2 }{ a^2 }=1\] in other words the \(a\) goes with the larger denominator since \(a^2=16\) we have \(a=4\) and since you know how it is oriented you have the vertices are \(4\) units above and below \((5,-3)\)

OpenStudy (anonymous):

you got those?

OpenStudy (anonymous):

would that be (5,1) (5,-7)?

OpenStudy (anonymous):

you sure?

OpenStudy (anonymous):

just kidding, it is right

OpenStudy (anonymous):

haha i was about to go and change my mind again :p

OpenStudy (anonymous):

ok now for the foci this time it is a little different we need \(c\) but here we use \(c^2=a^2-b^2\)

OpenStudy (anonymous):

that gives \(c^2=16-4=12\) so \(c=\sqrt{12}=2\sqrt3\)

OpenStudy (anonymous):

c^2=12?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but the square root of 12, so the answer would actually be 2 square root 3?

OpenStudy (anonymous):

again above and below \((5,-3)\) and again you just write it

OpenStudy (anonymous):

nevermind i forgot we have to find the square of 12!

OpenStudy (anonymous):

yeah it is what you said \(2\sqrt3\)

OpenStudy (anonymous):

so foci are \[(5,-3-2\sqrt3),(5,-3+2\sqrt3)\]

OpenStudy (anonymous):

how about the eccentricity?

OpenStudy (anonymous):

e=c/a, so would that be 2√3/4?

OpenStudy (anonymous):

yeah, but if you don't want your teacher to think you are challenged in the arithmetic department, you might want to cancel a 2 and write \(\frac{\sqrt3}{2}\)

OpenStudy (anonymous):

LOL good idea :D

OpenStudy (anonymous):

last part is nothing length of major axis is \(2a=8\) and length of minor axis is \(2b=4\) since \(a=4\) and \(b=2\)

OpenStudy (anonymous):

how do i find the length of the major and minor axis?

OpenStudy (anonymous):

see above

OpenStudy (anonymous):

oh ocay that is easy! so minor is 4 an

OpenStudy (anonymous):

and major is 8

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you SO much, i really appreciate all of your help!!!!

OpenStudy (anonymous):

awesome, thank you :)

OpenStudy (anonymous):

don't be confused by the "semimajor" axis etc that is just half of the major axis good luck on your exam

OpenStudy (anonymous):

thanks i truly appreciate it:)